if lim x tends to 0 (x^-3Sinx + ax^-2 + b) exists and is equal to zero then find value of a and b.. pls I need the way of finding b...

Grade:12

2 Answers

shiddhant bhattacharya
11 years ago

is the ans b=1/6 and a=R?

k sesi preetam
10 years ago

b=1/6 and a= -1